OpenStudy (anonymous):

Help me find the integral of (x*e^2x)/(1+2x)^2

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

you gota perform some log dances here, show me your working out

OpenStudy (anonymous):

I'm supposed to do it with integration by parts. I started off my considering (x*e^2x)=U and 1/(1+2x)^2 =dv but I'm not really getting anywhere with this

OpenStudy (anonymous):

where are you from btw? @ uni atm?

OpenStudy (anonymous):

Yeah I'm at Uni, this is from Calculus II. Finals are over but I'm trying to gain complete mastery before I move on.

OpenStudy (anonymous):

i'm in 11th grade

OpenStudy (anonymous):

anyway integral..hmm

OpenStudy (anonymous):

i'll write it out on paper first if that's okay.. :P a bit rusty

OpenStudy (anonymous):

What is Uni?

OpenStudy (anonymous):

@Pquevedo2 , \[\int\limits_{}^{} {{xe^{2x}dx} \over {(1+2x)^2}}\] Is this your question??

OpenStudy (anonymous):

that is my question, yes.

OpenStudy (anonymous):

I would begin with a u substitution (u=1+2x). That will simplify your denominator considerably, and break your integral into two pieces. Each piece should be solvable by integration by parts.

OpenStudy (anonymous):

Basically you will need to solve the INTEGRAL of e^u / u^2. You can do this by integration by parts, where you integrate the (1/u^2) piece and differentiate the (e^u) piece.

OpenStudy (anonymous):

rewrite it as: \[\int\limits [xe^(2x)]{dx/[1+2x^2]}\] let: [x e^(2x)] = u → {e^(2x) + x [2e^(2x)]} dx = du → {e^(2x) + 2x e^(2x)} dx = du → factoring out e^(2x) continuing...

OpenStudy (anonymous):

i had a tech glitch with the equation thingy lol

OpenStudy (anonymous):

[e^(2x)](1 + 2x) dx = du dx /[(2x + 1)^2] = dv → dividing and multiplying by 2, (1/2) {2dx /[(2x + 1)^2]} = dv → (1/2) {d(2x + 1) /[(2x + 1)^2]} = dv → (1/2) [(2x + 1)^(-2)] d(2x + 1) = dv → (1/2) [(2x + 1)^(-2+1)]/(-2+1) = v → (1/2) [(2x + 1)^(-1)]/(-1) = v → -1 /[2(2x + 1)] = v

OpenStudy (anonymous):

thus, integrating by parts, you get: ∫ u dv = u v - ∫ v du → ∫ [x e^(2x)] {dx /[(2x + 1)^2]} = [x e^(2x)]{-1 /[2(2x + 1)]} - ∫ {-1 /[2(2x + 1)]} [e^(2x)](1 + 2x) dx = (-1/2){[x e^(2x)] /(2x + 1)} + ∫ {[e^(2x)](1 + 2x) /[2(2x + 1)]} dx = canceling (2x + 1), (-1/2){[x e^(2x)] /(2x + 1)} + (1/2) ∫ e^(2x) dx = (-1/2){[x e^(2x)] /(2x + 1)} + (1/2) [(1/2)e^(2x)] + C = factoring out [(1/2)e^(2x)], [(1/2)e^(2x)] {- [x/(2x + 1)] + (1/2)} + C = [(1/2)e^(2x)] {(- 2x + 2x + 1) /[2(2x + 1)]} + C = [(1/2)e^(2x)] {1 /[2(2x + 1)]} + C thus, in conclusion: ∫ [x e^(2x)] / [(2x + 1)^2] dx = (1/4) {[e^(2x)] /(2x + 1)} + C

OpenStudy (anonymous):

medal please

OpenStudy (anonymous):

I would take t= 2x \[(1/4) \int\limits_{}^{}{ {te^{t} dt} \over {(1+t)}^2}\] Then add +1 and -1 \[(1/4) \int\limits_{}^{}{ {(t+1-1)e^{t} dt} \over {(1+t)}^2}\] \[(1/4) \int\limits_{}^{}{ {(t+1)e^{t} dt} \over {(1+t)}^2}- (1/4) \int\limits_{}^{}{ {e^{t} dt} \over {(1+t)}^2}\] \[(1/4) \int\limits_{}^{}{ {e^{t} dt} \over {(1+t)}}- (1/4) \int\limits_{}^{}{ {e^{t} dt} \over {(1+t)}^2}\] \[(1/4)\int\limits_{}^{} e^tdt( { {1} \over{ (1+t) } } - { {1} \over {1+t}^2}) \] Now use this property \[\int\limits_{}^{} e^xdx( f(x)+f'(x)) = e^x (f(x))+c\] In our case \[f(x)= {{1} \over {1+t}}\] \[f'(x)= {{-1} \over ({1+t})^2}\] Therefore we get \[(1/4)e^t ({{1} \over {1+t}})+C \] Now substitute t = 2x to get \[(1/4)e^{2x} ({{1} \over {1+2x}})+C \]

OpenStudy (anonymous):

wow i need to learn how to do that^

OpenStudy (anonymous):

i suck at using the Equation button :P

OpenStudy (anonymous):

No problem @Tai . What finally matters is the answer . :D Medal to you @Tai

OpenStudy (anonymous):

likewise!

OpenStudy (anonymous):

no medal for me :'(

OpenStudy (anonymous):

@tai Please explain these last three steps: [(1/2)e^(2x)] {- [x/(2x + 1)] + (1/2)} + C = [(1/2)e^(2x)] {(- 2x + 2x + 1) /[2(2x + 1)]} + C = [(1/2)e^(2x)] {1 /[2(2x + 1)]} + C How did you add the 1/2 to the - [x/(2x + 1)]? Thanks in advance